### v1

# Improving Temporal Understanding Logic Consistency in Video-Language Models via Attention Enhancement

Abstract
Large language models (LLMs) often generate self-contradictory outputs, which severely impacts their reliability and hinders their adoption in practical applications. In video-language models (Video-LLMs), this phenomenon recently draws the attention of researchers. Specifically, these models fail to provide logically consistent responses to rephrased questions based on their grounding outputs. However, the underlying causes of this phenomenon remain underexplored. In this work, we adopt an interpretability-driven approach to analyze, statistically summarize, and intervention the potential factors of the phenomenon. We find that one of the primary reasons for the inconsistency in responses lies in the inability of cross-modal attention heads to effectively distinguish video tokens across different timestamps. To address this, we propose an attention enhancement method called Temporally Conditioned Attention Sharpening (TCAS), which constructs an enhancement objective based on attention distinctions to enhance the model’s temporal resolution capability, thereby improving its temporal understanding logic consistency. Experimental results demonstrate that our method significantly enhances the temporal logic consistency of Video-LLMs. Further interpretability analyses reveal that our method indeed improves the temporal discriminability of attention heads, validating our conclusions. Additionally, our method achieves performance improvements in general video temporal grounding tasks, highlighting that temporal logic consistency is a bottleneck in temporal understanding. By enhancing consistency, our method drives significant progress in video temporal understanding.
Improving Temporal Understanding Logic Consistency in Video-Language Models via Attention Enhancement
Chengzhi Li   Heyan Huang   Ping Jianthanks:   Corresponding author   Zhen Yang   Yaning Tian School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China {lichengzhi,hhy63,pjian,bityangzhen,1120222204}@bit.edu.cn
1 Introduction
In recent years, Multimodal Large Language Models (MLLMs) have achieved significant strides in visual-language understanding Yue et al. (2024); Ying et al. (2024). Among them, Video Large Language Models (Video-LLMs) excel by combining video-based visual information with the pre-trained knowledge of large language models, enabling better comprehension and generation of natural language related to video content Tang et al. (2025). These models have demonstrated outstanding performance in tasks like video question answering and captioning, driving progress in multimodal intelligence Fu et al. (2024). However, the redundancy in video data challenges large language models in fine-grained temporal understanding. To address this, many studies aim to enhance the temporal understanding Guo et al. (2025); Ren et al. (2024), particularly for fine-grained tasks like Video Temporal Grounding (VTG).
These works often rely on additional specialized temporal modules to enhance Video-LLMs by providing additional temporal information streams Guo et al. (2025); Ren et al. (2024); Hu et al. (2024). Jung et al. Jung et al. (2024) proposed a benchmark to assess the consistency of state-of-the-art (SOTA) Video-LLMs, revealing their significant limitations in fine-grained temporal understanding. Specifically, Jung et al. Jung et al. (2024) find that all Video-LLMs struggle to provide logically consistent responses to rephrased questions, highlighting a lack of true temporal comprehension in both the models and the enhancement methods.
Although many modular methods have been proposed to enhance temporal understanding, the underlying reasons for large language models’ limitations in this area remain underexplored. Jung’s findings Jung et al. (2024) further reinforce the necessity for deeper analysis. Our work begins with temporal inconsistency, constructing an interpretability-driven analysis to uncover its intrinsic causes and deeply exploring temporal understanding. To our knowledge, this is the first study to analyze the temporal understanding consistency of Video-LLMs from an interpretability perspective. We focus on the following research questions:
1). What factors influence temporal understanding consistency in Video-LLMs, and how do these internal factors affect it? In Section 2, we use key head identifying, attention visualization, statistical summarization, and causal intervention methods, investigating the relationship between temporal understanding consistency and attention distributions. Finally, our findings reveal and confirm that the primary cause of response inconsistency lies in the inability of cross-modal attention heads to effectively differentiate video tokens across different timestamps.
2). How can we enhance the temporal understanding consistency of Video-LLMs based on the identified internal factors? In Section 3, we introduce Temporally Conditioned Attention Sharpening (TCAS), an optimization objective based on attention distributions. This compels the cross-modal attention heads to explicitly judge the relevance of temporally distributed information, improving the model’s temporal logic consistency. In Section 4, we validate the effectiveness across various experimental settings, demonstrating its ability to enhance temporal discriminability. Moreover, our method improves performance in general VTG tasks, highlighting temporal logic consistency as a key bottleneck.
Figure 1: Overview of our work. We first analyze to gain conclusions about the factors influencing temporal understanding logic consistency in Video-LLMs, and then propose a method to enhance the temporal consistency based on these conclusions.
The workflow is shown in Figure 1. The contributions of this paper are summarized as follows:
  • We reveal the critical role of minority key cross-modal attention heads in temporal understanding, particularly in maintaining temporal understanding logic consistency.
  • We conclude that the temporal discriminability of attention plays a crucial role in enhancing the model’s prediction of temporal logic consistency.
  • We propose TCAS, which improves temporal logic consistency by optimizing attention distributions, validated through extensive experiments.
2 Exploring Temporal Logic Consistency
We conduct experiments using the original grounding (G.), rephrased grounding (R-G.), and shifted grounding (S-G.) subsets of Charades-CON Jung et al. (2024). For each sample, we define the model’s Rephrase Consistency Score $c^{v}_{rg.}$ and Shift Consistency Score $c^{v}_{sg.}$ as $I^{v}_{ori}\times I^{v}_{rg.}$ and $I^{v}_{ori}\times I^{v}_{sg.}$, respectively, where $I^{v}_{ori}$, $I^{v}_{rg.}$, and $I^{v}_{sg.}$ represent the Intersection over Union (IoU) between the model predictions and the ground-truth timestamps for original, rephrased, and shifted grounding tasks.
This section focuses on analyzing the TimeChat-7b model Ren et al. (2024), which is a multimodal large language model used for long Video understanding and accurate time localization tasks, fusing temporal features and visual information through Sliding Video Q-Former and Timestation-aware frame encoder. In addition, for Generalizability of our conclusions, we analyze Qwen2.5-VL Bai et al. (2025) on the Event Order Judgment (EOJ) task in Section 2.4.
2.1 Detection and Interpretation
TimeChat utilizes a Q-Former module to transform video features extracted by the Encoder into visual tokens, which are then fed into the large language model. Therefore, we start from these visual tokens to locate and visualize the attention heads that particularly focus on visual tokens, and explore how their behavior affects temporal logic consistency performance. For any attention head $h$, its Cross-Modal Score on video sample $v$ is defined as the average of the attention scores from all event text tokens to visual tokens, specifically defined as follows:
$S^{h,v}_{cross.}=\frac{1}{|T|}\sum_{q\in T}\sum_{k\in V}A^{h,v}_{q,k}$ (1)
where $S^{h,v}_{cross.}$ denotes the cross-modal score, $T$ represents the tokens set of questions and answer text, $V$ represents the set of visual tokens, and $A^{h,v}_{q,k}$ is the attention score from text token as query token $q$ to visual token as key token $k$.
Figure 2: The distribution of cross-modal attention heads in TimeChat. The x-axis represents the attention layer index, and the y-axis represents the head index.
By sorting attention heads based on their cross-modal scores, we can identify those that particularly focus on visual tokens, which we refer to as cross-modal attention heads. The distribution of these cross-modal attention heads is shown in Figure 2. It is observed that only a small portion of mid-layer attention heads in the TimeChat model exhibit significantly higher cross-modal scores. These cross-modal attention heads align video information with text semantics, enabling the model to understand video content. How do these cross-modal attention heads establish the model’s temporal cognition?
To further investigate this question, we conduct a visual analysis of the behavior of cross-modal attention heads in the temporal dimension. The tokens in the key dimensions are aggregated based on their corresponding timestamps to highlight the attention head’s discriminability in the temporal dimension while disregarding spatial dimensions. In Figure 3, we present the distribution of attention scores for high-score cross-modal attention head $A^{14,3}$ across various samples.
Figure 3: Visualization of attention score distributions for key head $A^{14,3}$ across various samples. The start and end frames of the query event are marked in red at the corresponding token positions.
IoU 9MOH2 WIOOY KXIMH ZUU2X
$I_{ori}$ 0.956 0.451 0.413 0.414
$I_{sg.}$ 0.594 0.639 0 0
$c_{sg.}$ 0.562 0.288 0 0
Table 1: Prediction of samples for attention visualization.
Figure 3 shows how key head $A^{14,3}$ maps event text tokens to corresponding temporal visual tokens, enabling cross-modal alignment. However, $A^{14,3}$ fails to focus on the segments corresponding to the events in videos KXIMH and ZUU2X, correlating with poor consistency performance (Table 1). This suggests that this property of cross-modal attention heads could be a critical factor influencing the model’s temporal logic consistency. Additional patterns of other heads and samples can be found in Appendix A.1.
2.2 Statistical Analysis
We define Attention Discriminability Score as a measure of an attention head’s ability to differentiate event time segments within a video sample. This definition allows us to analyze the relationship between attention head behavior and the model’s consistency performance. Specifically, for any attention head $h$ and a video sample $v$ containing the query event $e$, attention discriminability is quantified using the ratio of the attention scores assigned to video tokens within the ground-truth time range to the total attention scores assigned to all video tokens. The formal definition is as follows:
$S^{h,v}_{disc.}=\frac{1}{|T_{e}|}\times\sum_{q\in T_{e}}\frac{\sum_{k\in V_{gt}}A^{h,v}_{q,k}}{\sum_{k\in V}A^{h,v}_{q,k}}$ (2)
$S^{v}_{disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{disc.}$ (3)
where $T_{e},V_{gt}$ represent the set of text tokens of event $e$ and the set of visual tokens in the ground-truth time range for event $e$, respectively. $S^{h,v}_{disc.}$ is the temporal discriminability score of attention head $h$ for video sample $v$, $H_{t}$ means the set of the top $t$ cross-modal score attention heads, and $S^{v}_{disc.}$ is the average temporal discriminability score of the top $t$ cross-modal attention heads for video sample $v$.
Figure 4: Violin plot showing the distribution of attention discriminability scores for different ranges of consistency scores. $\mu$, m, n denote the mean, median, and number of the discriminability score distribution, respectively.
For each video sample $v$, we calculate its consistency scores $c^{v}_{rg.}$, $c^{v}_{sg.}$ and attention discriminability score $S^{v}_{disc.}$. The Pearson correlation coefficient is computed by aggregating the values of these two metrics across all samples in the Charades-CON dataset. The discriminability scores distribution across different consistency sets is shown in Figure 4 as a violin plot, where we can find that $S_{disc.}$ and $c^{v}$ are positively correlated. The Pearson correlation coefficients are $0.4778$ and $0.4788$ for the rephrased grounding and the shifted grounding sets. This correlation coefficient is evidently statistically significant for Video-LLMs, which are inherently complex nonlinear systems. This result demonstrates a unambiguous positive correlation between the temporal discriminability of key heads and the model’s consistency performance. The scatter plot, the least-squares regression line, and more discussions are shown in Appendix A.2.
2.3 Causal Intervention
Statistical correlation alone is often insufficient to establish a causal relationship Pearl (2009). To further validate the causal relationship between attention discriminability and temporal logic consistency, a causal intervention Pearl (2009) study is conducted to determine whether improving attention discriminability leads to a rise in temporal logic consistency. Specifically, an intervention operation is designed that improves attention discriminability. By evaluating the model’s consistency performance after the intervention, we aim to confirm the causal relationship between attention discriminability and temporal logic consistency.
Specifically, the attention adjustment method is as follows: First, the multi-layer and multi-head attention distributions of the model during inference are obtained. Then, for the selected cross-modal attention heads with high attention scores, their query-to-key attention distributions are mixed with the ground-truth distribution within a specified intervention intensity $\alpha$. This process guides the model to focus more on the ground-truth temporal segments during inference. After the intervention, inference is performed again, and the changes in model outputs and attention distributions are compared to analyze the impact of the intervention on the model’s temporal understanding capabilities.
$\alpha$ $S_{disc.}$ G R-G. S-G.
R0.5 R0.7 $\mathcal{M}$ R0.5 R0.7 $\mathcal{M}$ R0.5 R0.7 $\mathcal{M}$
w/o 0.30 75.3 57.5 65.8 90.2 79.9 81.8 45.8 24.0 41.4
0.2 0.42 75.4 58.0 66.0 90.7 80.8 82.6 46.4 22.8 41.8
0.4 0.54 75.4 58.5 66.0 90.0 79.5 81.9 46.0 24.5 42.4
0.6 0.65 75.1 57.6 65.4 90.0 78.1 80.9 45.2 23.3 42.9
0.8 0.77 74.4 57.3 64.6 89.0 76.5 79.8 46.0 23.9 43.8
1.0 0.86 73.3 54.3 63.2 88.5 74.3 78.5 47.7 23.5 44.2
Table 2: Post-intervention scores with different $\alpha$ values. $\mathcal{M}$, R0.5, and R0.7 represent mIoU, recall at IoU thresholds of 0.5 and 0.7.
The post-intervention discriminability scores and consistency scores are presented in Table 2. As shown in the table, mild interventions effectively improve the model’s performance and consistency across the three subsets, while overly strong interventions lead to performance degradation. Additionally, relatively mild interventions are sufficient to enhance the model’s $c_{rg.}$, whereas improving $c_{sg.}$ requires stronger intervention intensity, indicating that the shifted grounding task poses greater challenges for the model in maintaining consistent outputs. These results demonstrate a clear causal relationship between attention discriminability and temporal logic consistency.
2.4 Generalizability Analysis
To validate the generalizability of our conclusions across other models and tasks, we extend the analysis to the Qwen2.5-VL Bai et al. (2025). Furthermore, we select the Event Order Judgment (EOJ) task as a representative one of complex temporal reasoning tasks Yu et al. (2023); Swetha et al. (2025), which involves temporal order judgment, allowing us to further investigate the model’s temporal capabilities in complex multi-event scenarios.
Figure 5: Violin plot of analysis results on EOJ task.
In the EOJ task, we define a new consistency score and an attention discriminability score to verify our findings. The distribution of attention discriminability scores for different ranges of consistency scores is shown as a violin plot in Figure 5. We observe a significant positive correlation (Pearson correlation coefficient is $0.447$), consistent with the trend found in TimeChat. This demonstrates that the relationship between attention discriminability and temporal logic consistency holds true for other leading Video-LLMs. The detailed implementation can be found in Appendix A.3.
3 Attention Enhancement
In the previous section, it is confirmed that the temporal discriminability of cross-modal attention heads is a key factor influencing the model’s temporal logic consistency. How can we enhance attention discriminability to improve temporal consistency? To achieve generalization and cross-architecture applicability, we formulate the attention discriminability enhancement as an optimization objective termed Temporally Conditioned Attention Sharpening Loss (TCAS Loss). The specific loss computation is detailed in Algorithm 1, where the $\textsc{SelectTopHeads}(H,S,t)$ means selecting the top $t$ attention heads from the set of all heads $H$ based on scores $S$.
Algorithm 1 TCAS Loss
Input: Video and text tokens $V,T$, model $M$
Parameter: Top heads number $t$, margin $m$, threshold thrthr
Output: TCAS loss $\mathcal{L}$
1: Response $r$, attentions $A^{H}\leftarrow M(V,T)$.
2: Compute $S^{v}_{cross.}$ using Equation (1).
[…]
4:for each head $h$ in $H_{t}$ do
5:  Aggregate visual tokens in $A^{h}_{T,V}$ according to timestamps to get $A^{h}_{T,V_{time}}$
6:  $T^{h}_{valid}\leftarrow\{q\in T\mid\text{Max}(A^{h}_{q,V_{time}})>thr\}$
7:  for each text query token $q$ in $T^{h}_{valid}$ do
8:   Mean $M^{h}_{q}=\sum_{k\in V_{time}}(A^{h}_{q,k})/|V_{time}|$
9:   Pos. $P^{h}_{q}\leftarrow\{A^{h}_{q,k}\in A^{h}_{q,V_{time}}\mid A^{h}_{q,k}>M^{h}_{q}\}$
10:   Neg. $N^{h}_{q}\leftarrow\{A^{h}_{q,k}\in A^{h}_{q,V_{time}}\mid A^{h}_{q,k}<M^{h}_{q}\}$
11:   Loss $\mathcal{L}^{h}_{q}=\text{Max}(m+\text{Max}(N^{h}_{q})-\text{Min}(P^{h}_{q}),0)$
12:  end for
13:end for
14:return $(\sum\limits_{h\in H_{t}}\sum_{q\in T^{h}_{valid}}\mathcal{L}^{h}_{q})/(\sum\limits_{h\in H_{t}}|T^{h}_{valid}|)$
We first identify the attention heads with the highest cross-modal scores and the text tokens that focus on specific video timestamps. This approach preserves the original contextual interaction capabilities of the attention heads. A contrastive learning loss is designed for the selected text query tokens to amplify the differences in attention scores across distinct timestamps, thereby compelling the attention heads to make explicit relevance judgments for temporally distributed information. We use weight $w_{ae}$ to balance the TCAS loss and next token prediction loss. Our method introduces no additional modules, ensuring compatibility with various transformer-based architectures Vaswani et al. (2023). Furthermore, it does not depend on specific text templates or video patterns, making it broadly applicable to diverse video-language tasks.
Method Data FT Charades-CON ActivityNet-CON
Ground R-Ground S-Ground H-Verify C-Verify Ground R-Ground S-Ground H-Verify C-Verify
Qwen2.5vl VTune SFT 28.3 17.5 (62.0) 6.0 (21.1) 15.1 (53.3) 14.8 (52.1) 18.2 14.5 (79.8) 2.5 (13.5) 12.0 (66.0) 13.2 (72.6)
Qwen2.5vl VTune TCAS 34.0 23.0 (67.5) 8.1 (23.7) 19.6 (57.6) 18.5 (54.3) 18.7 15.1 (80.5) 2.5 (13.6) 12.3 (65.5) 14.0 (74.6)
videollama VTune SFT 54.4 38.2 (70.3) 10.9 (20.0) 30.7 (56.5) 30.0 (55.2) 33.0 24.7 (74.8) 10.0 (30.2) 20.2 (61.1) 17.7 (53.7)
videollama VTune TCAS 62.4 44.4 (71.2) 13.4 (21.4) 34.8 (55.8) 37.3 (59.7) 33.8 25.2 (74.7) 11.1 (32.7) 20.9 (61.8) 18.6 (55.1)
TimeChat 30.5 25.0 (82.1) 5.6 (18.5) 14.0 (45.9) 15.6 (51.2) 4.6 2.9 (64.1) 1.0 (21.2) 2.1 (46.7) 2.4 (52.2)
TimeChat TimeIT SFT 55.8 50.9 (91.3) 10.5 (18.9) 16.7 (30.0) 25.7 (46.2) 25.3 20.2 (80.4) 7.5 (29.9) 8.7 (34.5) 12.6 (49.9)
TimeChat TimeIT TCAS 66.2 60.7 (91.3) 33.8 (51.1) 37.3 (56.3) 31.4 (47.4) 38.1 29.2 (76.6) 10.4 (27.3) 24.2 (63.3) 18.3 (47.9)
TimeChat VTune SFT 76.2 69.2 (90.8) 36.2 (47.5) 44.8 (58.8) 42.4 (55.7) 37.4 28.3 (75.6) 10.6 (28.3) 19.6 (52.3) 19.3 (51.5)
TimeChat VTune TCAS 83.3 75.0 (90.1) 39.5 (47.4) 52.9 (63.5) 50.8 (61.0) 38.8 31.0 (79.9) 11.1 (28.6) 20.8 (53.6) 21.5 (55.5)
Table 3: Comparison results for consistency and grounding performance. For the grounding task, IoU scores are reported, while accuracy rates are provided for the Verify task. Each cell shows the absolute performance, with relative performance in parentheses. Qwen2.5vl and videollama refers to Qwen2.5-VL-7B-Instruct and Video-LLaMA-7B.
4 Experiments
In this section, we present a series of experiments to evaluate the effectiveness of our proposed method in enhancing grounding performance and logic consistency in the VTG task. This section describes the experimental setup, presents comparison results, analyzes component contributions, provides some prediction examples, and reports the attention discriminability scores after enhancement.
4.1 Experimental Setup
4.1.1 Models
We conduct experiments on the Qwen2.5-VL Bai et al. (2025), Video-LLaMA Zhang et al. (2023), and TimeChat Ren et al. (2024) models, which respectively represent Video-LLMs based on MLP projection, Video-LLMs based on Qformer, and time-aware models. More details about these models can be found in Appendix B.1.
4.1.2 Datasets
We conduct experiments on two widely-used VTG datasets: ActivityNet Captions Krishna et al. (2017) and Charades Gao et al. (2017). ActivityNet Captions comprises approximately 20k untrimmed YouTube videos with an average duration of 120 seconds, while Charades contains 9848 short videos with an average length of 30 seconds. Additionally, Charades-CON and ActivityNet-CON, introduced by Jung et al. Jung et al. (2024), serve as consistency benchmarks to ensure fair, transparent, and reliable comparisons.
4.1.3 Tuning Annotations
We select two different training Annotations. The first is a standard instruct annotation dataset for the VTG consistency task, called TimeIT, proposed by Ren et al. Ren et al. (2024). The second dataset is VTune, introduced by Jung et al. Jung et al. (2024), which has been shown to effectively enhance the consistency of Video-LLMs. Our method demonstrates improved results over baseline fine-tuning methods using both datasets, indicating that it does not rely on specific data sources. The dataset examples can be found in Appendix B.2.
4.1.4 Evaluation Metrics
Following previous works Jung et al. (2024); Ren et al. (2024), we evaluate grounding performance using IoU metrics, while measuring the consistency of the model responses through the accuracy rate. Details about the evaluation metrics can be found in Appendix B.3.
4.1.5 Implementation Details
We implement our method using PyTorch on a single NVIDIA A100 GPU. All settings are trained with the Adam optimizer (learning rate: $1\times 10^{-5}$, batch size: 4) for approximately 3 days. Qwen2.5-VL, TimeChat, and Video-LLaMA process 16, 96, and 8 frames respectively, all at $224\times 224$ spatial resolution.
Method Charades-STA ActivityNet-Cap.
R@1,0.5 R@1,0.7 R@1,0.5 R@1,0.7
Task-Specific Models
BM-DETR Jung et al. (2025) 59.4 38.3 49.6 30.6
Mr.BLIP Meinardus et al. (2024) 69.3 49.2 53.9 35.5
Video-LLMs
HawkEye Wang et al. (2024) 58.3 28.8 34.7 17.7
VTG-LLM Guo et al. (2025) 57.2 33.4 - -
videollama (w. SFT) 37.1 20.1 34.3 19.1
videollama (w. TCAS) 37.5 20.1 35.2 20.0
Qwen2.5vl (w. SFT) 27.3 14.1 16.5 7.8
Qwen2.5vl (w. TCAS) 32.4 16.6 16.5 7.9
TimeChat Ren et al. (2024) 46.7 23.7 28.0 15.8
TimeChat (w. SFT) 58.4 34.7 41.0 23.7
TimeChat (w. TCAS) 60.2 37.6 41.2 24.9
Table 4: Comparison results for VTG performance on Charades-STA and ActivityNet-Caption. R@1,0.5 and R@1,0.7 represent Recall at IoU thresholds of 0.5 and 0.7.
4.2 Comparison Experiments
We conducted experiments to evaluate our method’s grounding performance and consistency, demonstrating the effectiveness of our method, TCAS, in improving temporal logic consistency and understanding.
4.2.1 Baselines
We evaluate our method’s effectiveness by comparing it with several baseline models and approaches. For open-source models, direct comparison without fine-tuning would be unfair, as no existing modeling methods specifically target consistency enhancement. Therefore, we compare our approach with standard fine-tuning methods, specifically Supervised Fine-Tuning (SFT) Zhang et al. (2024), applied to several selected backbones. The results of other models without tuning include GPT-4o OpenAI (2024), Gemini-1.5 Flash Team (2024), and so on, listed in Appendix B.4.
Notably, some models, such as VTG-LLM Guo et al. (2025), focus on improving temporal understanding without addressing consistency issues. So we only compare their grounding performance, excluding consistency metrics.
4.2.2 Main Results
We present the comparison results on consistency (Table 3) and grounding performance (Table 4). The best results are in bold or underlined in different areas. Light gray rows indicate results from our method, TCAS.
4.2.3 Discussion
We focus on the following phenomena:
1). The improvements brought by TCAS are evident across different models, training datasets, and benchmarks. As shown in Table 3, TCAS achieves favorable results in all comparisons, indicating that TCAS is not dependent on specific models, benchmarks, or training annotations and can robustly enhance model performance and consistency. The performance improvement on ActivityNet-CON is relatively smaller, which is due to the noisier data distribution of this dataset, where event descriptions are presented in long sentences that deviate from the general understanding of the concept of events. Consequently, all methods and models perform worse on this dataset, making the smaller improvement acceptable.
2). TCAS improves both grounding and consistency performance. On the classic VTG benchmarks, Charades-STA and ActivityNet-Captions, TCAS achieves SOTA performance among Video-LLMs methods. This suggests that inconsistent temporal understanding is a key limitation for the grounding performance of these models. By enhancing attention discriminability, TCAS improves temporal logic consistency, thereby boosting grounding performance and advancing video temporal understanding.
Hyperparameters Charades-CON
t{t} $m$ thrthr $w_{ae}$ G. R-G. S-G. H-V. C-V.
Analysis of top heads number $t$
16 0.2 0.1 0.5 80.91 72.14 36.77 40.98 42.17
48 0.2 0.1 0.5 77.37 69.66 39.04 45.04 44.24
Analysis of margin $m$
32 0.15 0.1 0.5 80.76 71.61 36.59 51.23 48.26
32 0.25 0.1 0.5 78.64 71.07 34.35 49.48 45.09
Analysis of threshold thrthr
32 0.2 0.05 0.5 81.90 74.95 41.45 54.30 48.95
32 0.2 0.15 0.5 81.05 72.33 34.84 44.10 40.65
Analysis of loss weight $w_{ae}$
32 0.2 0.1 0.3 80.91 74.76 39.04 48.67 48.75
32 0.2 0.1 0.7 80.48 71.10 35.55 48.69 46.59
Empirical optimal combination
32 0.2 0.1 0.5 83.31 75.02 39.52 52.93 50.81
Table 5: Ablation study results for hyperparameter sensitivity analysis. The table shows the absolute consistency scores for different hyperparameter configurations.
4.3 Ablation Study
Figure 6: Responses of TimeChat on grounding and verification for sample EQ4GN before and after TCAS enhancement.
Our method is not composed of distinct independent modules, making traditional ablation studies infeasible. However, it depends on several key hyperparameters (i.e., top heads number $t$, margin $m$, threshold thrthr, and loss weight $w_{ae}$) that regulate the intensity of attention adjustment from various perspectives. Sensitivity analysis of these hyperparameters allows us to assess their contributions to improving temporal logic consistency and to further examine the intrinsic effectiveness of our approach.
The impact of different hyperparameters on model performance varies. The parameters $t$ and thrthr primarily determine the adjustment scope of TCAS on the aspect of heads and tokens, respectively, while $m$ and $w_{ae}$ control the intensity of TCAS enhancement. As shown in Table 5, increasing $t$ leads to a significant drop in G. and R-G. subsets, while a larger thrthr results in a substantial decline in S-G., H-V., and C-V. subsets, indicating that these two parameters are most critical for model performance. Overall, our method is more sensitive to scope-related hyperparameters while maintaining robustness to intensity-related hyperparameters.
4.4 Case Study
To visually demonstrate the effectiveness of our method, we select several prediction results from Charades-CON Jung et al. (2024) for a case study. For the example shown in Figure 6, our model successfully grounds query events in the video and correctly answers the questions, demonstrating improved temporal reasoning capabilities compared to baseline methods. This result highlights the potential of our approach to enhance the understanding of temporal dynamics in video content. More examples are provided in the Appendix B.5.
4.5 Statistical Analysis
To confirm that TCAS improves consistency by enhancing attention discriminability, rather than through unintended artifacts, we analyze the distribution of the top-32 attention discriminability scores for both the baseline and TCAS-enhanced models on the Charades-CON dataset. As depicted in the histogram in Figure 7, the statistical analysis reveals a distinct rightward shift in the score distribution after enhancement. This demonstrates that TCAS effectively boosts attention discriminability, which consequently improves consistency, thereby reinforcing the validity of our conclusions.
Figure 7: Discriminability scores distribution 'Before' (baseline model) and 'After' (TCAS-enhanced model) attention enhancement. The KDE means Kernel Density Estimation curves.
5 Related Work
5.1 Video-LLMs on VTG
Recent Video-LLMs excel in fine-grained temporal understanding tasks like VTG by independently encoding timestamps and interacting with video information to capture precise temporal segments, marking a significant improvement over traditional VTG methods by enabling diverse natural language outputs Guo et al. (2025); Ren et al. (2024); Hu et al. (2024). New benchmarks for fine-grained temporal understanding have been proposed to analyze their limitations Fu et al. (2024); Jung et al. (2024). For example, the consistency benchmark by Jung et al. Jung et al. (2024) highlights response inconsistency. Unlike prior work, we focus on the internal representations of Video-LLMs rather than solely developing new benchmarks for evaluation, providing deeper insights into their temporal cognition.
5.2 Interpretability Analysis in MLLMs
In recent years, interpretability of LLMs has emerged as a key research focus to enhance model transparency and controllability Rai et al. (2025). However, the understanding of MLLMs remains limited Lin et al. (2025). Nikankin Nikankin et al. (2025) and Venhoff Venhoff et al. (2025) investigated MLLMs’ image understanding using circuit and SAE, respectively. Li et al. Li et al. (2025) probed visual features and identified the LLM decoder as the bottleneck for temporal reasoning, but did not explore the root causes of this limitation. From a consistency perspective, we offer a deeper explanation, attributing the decoder’s limitations to the insufficient discriminability of cross-modal attention heads.
6 Conclusion
In this paper, we first analyze the role of cross-modal attention heads in establishing temporal cognition and then validate the causal relationship between attention discriminability and temporal logic consistency through correlation analysis and causal intervention.
Based on these insights, we introduce a novel method, TCAS, that optimizes attention distributions to improve temporal discriminability. Our experimental results demonstrate that our method effectively enhances temporal logic consistency in Video-LLMs, achieving SOTA grounding performance and consistency performance across various settings, while other extensive experiments further demonstrate its reliability.
Appendix A More Details about Analysis
A.1 More Attention Patterns
Figure 8: Visualization of attention score distributions for key heads across various samples. Subplots a-d are high discriminability cases, while e and f represent cases where the key heads fail to distinguish correctly.
In the main text, we presented the attention patterns of key attention heads in TimeChat. To further support the analysis in Section 2.1, we provide additional attention patterns of key attention heads in TimeChat. As shown in Figure 8, we observe that the attention patterns of these heads are similar to those shown in the main text, with a focus on the temporal dimension. This indicates that these heads are indeed crucial for understanding the temporal dynamics of video content.
A.2 Statistical Analysis Details
Figure 9: Scatter plot showing the relationship between the consistency score and the attention discriminability score, with a least-squares regression line included.
In the main text, we conduct the correlation analysis of attention discriminability scores and consistency scores in TimeChat. To further support the analysis in Section 2.1, we provide the scatter plot with the least-squares regression line to illustrate the relationship between these two metrics.
As shown in Figure 9, we observe that as the consistency score increases, the distribution of attention discriminability scores shifts to the right. The points clustered on the left side of the x-axis in the figure correspond to difficult samples where the model’s predicted IoU is 0.
For rephrased grounding and shifted grounding, the Pearson correlation coefficients between the two metrics are $0.4778$ and $0.4788$, respectively, with p-values of $\mathbf{5.18\times 10^{-41}}$ and $\mathbf{3.69\times 10^{-41}}$, both far below $\mathbf{0.05}$, indicating a statistically significant positive correlation between the two metrics. Since Video-LLMs are inherently complex nonlinear systems, the figure shows that the two metrics are not linearly correlated. However, their positive correlation is statistically significant, as evidenced by the p-values.
A.3 Details of Analysis on EOJ Task
To explore the correlation between attention discriminability and consistency performance on the Event Order Judgement (EOJ) task, we constructed a probing dataset based on the annotations from Charades-STA test subset, containing 1,000 consistent question-answer pairs. Each pair includes a video clip and multiple logically equivalent questions that require the model to determine the order of different events in the video clip. An example of the question-answer pairs is shown in Figure 10.
Figure 10: An example of the EOJ Probing dataset. Each sample contains two events with a clear order, and four sets of eight logically equivalent questions are constructed for their order relationship.
A.3.1 Detection and Interpretation.
Using the same method as in the main text, we first identify the cross-modal attention heads in Qwen2.5-VL that are relevant to the EOJ task. The distribution of these heads is shown in Figure 11. We observe that, consistent with the analysis in the main text, the cross-modal scores in Qwen2.5-VL are also primarily concentrated in a few heads in the middle layers.
Figure 11: The distribution of cross-modal attention heads in Qwen2.5-VL. The x-axis represents the attention layer index, and the y-axis represents the head index.
Next, we visualize the attention patterns of key heads on several samples, as shown in Figure 12. We find that the attention patterns of these heads are consistent with the analysis in the main text, establishing a mapping relationship between the query tokens of event queries and the video key tokens corresponding to the time range of events through attention scores, thus achieving cross-modal temporal relationship alignment.
Figure 12: Attention patterns of key heads in the EOJ task.
A.3.2 Correlation Analysis.
In the EOJ task, the consistency score $c^{v}$ is defined as follows:
$c_{v}=\frac{\sum_{q\in Q_{v}}F1(r_{v,q})}{|Q_{v}|},\quad v\in\mathcal{D}$ (4)
where $c_{v}$ indicates the model’s Consistency Score for video sample $v$, $r_{v,q}$ denotes the model’s response to question $q$ for video sample $v$, $Q_{v}$ refers to the set of questions associated with video sample $v$, and $\mathcal{D}$ represents the dataset of all video samples.
The Attention Discriminability Score $S^{h,v}_{disc.}$ is defined as the Kullback-Leibler divergence Kullback and Leibler (1951) between the average attention score distributions of two events, specifically:
$P^{h,v}_{e}=\frac{\sum_{q\in T_{e}}A^{h,v}_{q,V}}{\sum_{k\in V}\sum_{q\in T_{e}}A^{h,v}_{q,k}}$ (5)
$S^{h,v}_{disc.}=P^{h,v}_{e_{1}}\log\frac{P^{h,v}_{e_{1}}}{P^{h,v}_{e_{2}}}+P^{h,v}_{e_{2}}\log\frac{P^{h,v}_{e_{2}}}{P^{h,v}_{e_{1}}}$ (6)
$S^{v}_{disc.}=\frac{1}{|H_{t}|}\sum_{h\in H_{t}}S^{h,v}_{disc.}$ (7)
where $P^{h,v}_{e}$ denotes the average attention score distribution of all text tokens of event $e$ to the set of all visual tokens $V$, $T_{e}$ represents the set of text tokens of event $e$, $A^{h,v}_{q,V}$ is the attention score from text token as query token $q$ to all visual tokens $V$ as key tokens, $S^{h,v}_{disc.}$ is the temporal discriminability score of attention head $h$ for video sample $v$ containing two events $e_{1}$ and $e_{2}$, $H_{t}$ means the set of the top $t$ cross-modal score attention heads, and $S^{v}_{disc.}$ is the average temporal discriminability score of the top $t$ cross-modal attention heads for video sample $v$.
The correlation analysis of attention discriminability scores and consistency scores is shown in Figure 13 as a scatter plot with the least-squares regression line. We observe that as the consistency score increases, the attention discriminability score also increases significantly, indicating a clear positive correlation between attention discriminability and consistency performance. The corresponding violin plot is shown in Figure 5 of the main text.
Figure 13: Scatter plot showing the relationship between the consistency score and the attention discriminability score on EOJ task, with a least-squares regression line included.
In conclusion, our analysis is validated across both the Qwen2.5-VL model and the EOJ task, underscoring the robustness and generalizability of our conclusions. This provides compelling evidence that attention temporal discriminability is a key factor of performance for Video-LLMs in general temporal understanding tasks.
Appendix B More Details about Experiments
B.1 Backbone Models Introduction
Qwen2.5-VL Bai et al. (2025) is a representative Video-LLMs featuring a standard cross-modal architecture. We select Qwen2.5-VL to further validate our findings from the previous section. Video-LLaMA Zhang et al. (2023) is a classic video-LLM using Qformer and video Qformer for video-text alignment. TimeChat Ren et al. (2024), a SOTA video-LLM for the VTG task, integrates timestamp information into its architecture, enhancing its ability to understand the temporal dynamics of video content. We include Video-LLaMA and TimeChat to ensure a fair comparison with the current SOTA method.
B.2 Datasets Introduction and Examples
TimeIT
TimeIT is a dataset proposed by Ren et al. Ren et al. (2024) for the various temporal understanding tasks including description, video grounding and temporal reasoning. The splits Charades-VTG and ActivityNet-VTG of it is used in our experiments. These two splits contain 9,848 video clips with 12,408 question-answer pairs and 14,420 video clips with 35,692 question-answer pairs respectively.
VTune
VTune is a dataset reported as a method in the original paper Jung et al. (2024). By constructing event verification queries and temporal verification queries annotation data, it effectively enhances the consistency of Video-LLMs from a view of synthetic data augmentation. Two splits Charades-VTune and ActivityNet-VTune contains 9,848 video clips with 99,244 question-answer pairs and 14,420 video clips with 205,510 question-answer pairs.
CON
Charades-CON and ActivityNet-CON are two benchmarks proposed by Jung et al. Jung et al. (2024) to evaluate the consistency of Video-LLMs in temporal understanding tasks. These benchmarks contain 500 video clips with 707 question-answer pairs and 1,422 question-answer pairs.
The examples of TimeIT and VTune are shown in Figure 15. These two datasets are used for training, while the last two datasets are used for evaluation in our experiments.
Figure 14: Responses of TimeChat for sample 9T11N (Top) and T0HLX (Bottom) before and after TCAS enhancement.
Figure 15: Examples from the training datasets (TimeIT and VTune) used in our training experiments.
B.3 Evaluation Metrics
Following the previous works Jung et al. (2024), we use these metrics to evaluate the performance of our method:
  • IoU: The ratio of the intersection over the union between the predicted time range and the ground-truth time range.
  • Accuracy: The percentage of questions that can be answered correctly by the model. We use it evaluate absolute performance on the verification tasks (H-V. and C-V.) of CON datasets.
  • R@1,IoU=0.5: The percentage of the IoU of prediction is greater than 0.5. We use it to evaluate the absolute performance of the model on all grounding tasks.
  • R@1,IoU=0.7: The percentage of the IoU of prediction is greater than 0.7. We use it to evaluate the absolute performance of the model on all grounding tasks.
  • mIoU: The mean IoU of all questions in the dataset.
  • Consistency: All Consistency metrics are defined as the original grounding R@1,IoU=0.5 times corresponding absolute performance metrics.
B.4 Other Baselines
We also compare our method with the following closed-source models: GPT-4o OpenAI (2024) Gemini-1.5 Flash Team (2024) in Table 6. It is worth noting that these results are reported in Jung et al. (2024) and are listed here for reference only.
Models Consistency Metrics
G R-Ground S-Ground H-Verify C-Verify
Charades-CON
GPT-4o 28.5 21.2 (74.3) 9.3 (32.8) 17.8 (62.4) 20.3 (71.3)
gemini-1.5 34.6 29.7 (85.7) 24.8 (71.7) 22.8 (65.8) 24.5 (70.8)
ActivityNet-CON
GPT-4o 26.8 18.1 (67.5) 10.4 (38.8) 16.5 (61.7) 18.4 (68.8)
gemini-1.5 37.8 30.8 (81.4) 24.8 (65.6) 22.4 (59.3) 26.8 (70.8)
Table 6: Results of other models without tuning, listed for reference only Jung et al. (2024).
B.5 Case Study
In this section, we present more prediction examples in Figure 14 to illustrate the effectiveness of our method in enhancing temporal consistency in Video-LLMs. Our model successfully grounds key events and answers questions about the video, demonstrating improved temporal capabilities and consistency compared to baseline models.